OpenStudy (grandmoffchris):

7. -2x = -15 8. 6x = 42 9. g/3 = -0.7 10. y/4 = 3 11. -7 = j/2 12. y/-2 = -6

OpenStudy (grandmoffchris):

Will fan and medal Pls

OpenStudy (grandmoffchris):

@dan815 @AloneS @Abstract_Undercover @Englishguy SOMEONE HELP!!

OpenStudy (grandmoffchris):

K

OpenStudy (grandmoffchris):

Yes Pls

OpenStudy (grandmoffchris):

Thx

OpenStudy (grandmoffchris):

@Englishguy Didnt work, Didnt understand a thing.

OpenStudy (grandmoffchris):

Doesnt explain how to solve it

OpenStudy (grandmoffchris):

Just Help me!!!, have not so much time, DONT EXPLAIN JUST ANSWRR

OpenStudy (anonymous):

Either divide or multiply the number connectected to the variable to the number on the other side. for example 24=6y y/3=4 y=3 y=12

OpenStudy (grandmoffchris):

So that means what?

OpenStudy (anonymous):

open study no cheating policy

OpenStudy (grandmoffchris):

waait what

OpenStudy (grandmoffchris):

ur not helping.

OpenStudy (grandmoffchris):

@TheSmartOne @aaronq @dan815 Pls help, Im bad at this and i have not so much time, juts like 30 mins, Pls help.

OpenStudy (grandmoffchris):

JUST HELP ME

OpenStudy (grandmoffchris):

@sky060902 Help

OpenStudy (grandmoffchris):

someone???

pooja195 (pooja195):

-2x = -15 We are isolating x correct? Divide both sides by -2

OpenStudy (grandmoffchris):

Have no idea

OpenStudy (grandmoffchris):

Idk, I have millions of times, I... AM.... BAD.... AT.... THIS!!!

OpenStudy (grandmoffchris):

said that

OpenStudy (grandmoffchris):

@faithpeterson2021

OpenStudy (anonymous):

I don't know the answer, sorry but, I found a website you can use for math help http://www.freemathhelp.com/calculators.html

pooja195 (pooja195):

Our expression is: \[\huge~\rm~ -2x = -15\] When we want to isolate the expression we divide both ides by -2x because -2x/-2x/1 so you have x by it self -15/-2=15/2 - divided by a negative will result in a positive solution. so we have x=15/2 Also no one will want to help you if you don't participate instead of idk try to attempt the problem so that you will learn. I know i have said this to you before.

OpenStudy (grandmoffchris):

@girlmeetsworld

pooja195 (pooja195):

And to all the other users: Please don't give links to calculators. The user needs to learn. @GrandMoffChris quit tagging people.

OpenStudy (girlmeetsworld):

Pooja got this ... she better at math than me. P

OpenStudy (anonymous):

Yeah pooja is good at math u explained that perfectly

OpenStudy (nathalie0918):

:)

pooja195 (pooja195):

I want you to try 6x = 42 DO NOT say "idk"

OpenStudy (grandmoffchris):

Idk

pooja195 (pooja195):

Then i'm sorry i can't help you.

OpenStudy (grandmoffchris):

wait

OpenStudy (grandmoffchris):

7.5?

pooja195 (pooja195):

We are dividing BOTH sides by 6 6/6=x x=42/6 42 divided by 6=?

OpenStudy (grandmoffchris):

7

pooja195 (pooja195):

Congrats you solved it.

OpenStudy (grandmoffchris):

:D

pooja195 (pooja195):

Now this one g/3 = -0.7 When we are working with these we need to multiply by the reciprocal the reciprocal basically means the flipped fraction g/3=3/g \[\huge~\rm~\frac{g }{3}\times \frac{ 3 }{ g }=g\] The numbers cancel out and you are left with just g Now multiply \[\huge~\rm~\bf -0.7 \times 3=? \]

OpenStudy (grandmoffchris):

mmgggg

pooja195 (pooja195):

Tell me "idk" and i leave

OpenStudy (grandmoffchris):

i wont

OpenStudy (grandmoffchris):

-2.1?

pooja195 (pooja195):

Congrats you solved another one

pooja195 (pooja195):

Do this one using the same method y/4 = 3

OpenStudy (grandmoffchris):

Alright

OpenStudy (grandmoffchris):

12

pooja195 (pooja195):

Good :) y=12

OpenStudy (grandmoffchris):

:D

pooja195 (pooja195):

Next one -7 = j/2

OpenStudy (grandmoffchris):

mhhh

OpenStudy (grandmoffchris):

-14?

pooja195 (pooja195):

Well done :D Remeber to write j=-14

OpenStudy (grandmoffchris):

Ok

pooja195 (pooja195):

y/-2 = -6

OpenStudy (grandmoffchris):

Final one

pooja195 (pooja195):

mhm

OpenStudy (grandmoffchris):

12?

OpenStudy (grandmoffchris):

although there is one more

OpenStudy (grandmoffchris):

2/3m = -8

pooja195 (pooja195):

Same thing flip the fraction 2/3 x 3/2 = 1 m=3/2 x -2 =?

OpenStudy (grandmoffchris):

-3?

pooja195 (pooja195):

My bad i meant 3/2 x -8

pooja195 (pooja195):

Try again

OpenStudy (grandmoffchris):

oH

OpenStudy (grandmoffchris):

-12?

pooja195 (pooja195):

Well done

OpenStudy (grandmoffchris):

:D

pooja195 (pooja195):

=)

OpenStudy (grandmoffchris):

:DDDD

OpenStudy (grandmoffchris):

I got some more

OpenStudy (grandmoffchris):

-8 = b/-3 @pooja195

pooja195 (pooja195):

Try it yourself

OpenStudy (grandmoffchris):

Help

OpenStudy (grandmoffchris):

and pooja, still need to work it out with more training

OpenStudy (grandmoffchris):

or practice

OpenStudy (grandmoffchris):

So help me?

pooja195 (pooja195):

Flip and multiply,

OpenStudy (grandmoffchris):

Ok

OpenStudy (grandmoffchris):

-3 x -8?

pooja195 (pooja195):

yep

OpenStudy (grandmoffchris):

:D

pooja195 (pooja195):

:)

OpenStudy (grandmoffchris):

24?

OpenStudy (grandmoffchris):

Now i get it, :D, Thanks

OpenStudy (grandmoffchris):

I flip the numbers then multiply and thats all

OpenStudy (grandmoffchris):

Thank you so much

pooja195 (pooja195):

Yw ^_^

OpenStudy (grandmoffchris):

:D

OpenStudy (grandmoffchris):

@pooja195 i need help with one that there is 3 numbers, 3/8t = -27

pooja195 (pooja195):

Flip and multiply

OpenStudy (grandmoffchris):

wait but do i divide 3 / 8? and then multiply it?

pooja195 (pooja195):

Flipe 3/8=8/3 3/8 x 8/3= t t=8/3 x -27

OpenStudy (grandmoffchris):

Oh okay

OpenStudy (grandmoffchris):

-72?

pooja195 (pooja195):

Well done